OpenStudy (anonymous):

Solve A=h (a+b)for b a. b= (A-ha)/h b. b= (A-a)/h c. b= (A+ha)/h d. b= (A+a)/h

OpenStudy (anonymous):

@tyler**collins

OpenStudy (jordanloveangel):

Solution for a=h(a+b)forb equation: Simplifying a = h(a + b) * forb Reorder the terms for easier multiplication: a = h * bfor(a + b) Multiply h * bfor a = bfhor(a + b) a = (a * bfhor + b * bfhor) a = (abfhor + b2fhor) Solving a = abfhor + b2fhor Solving for variable 'a'. Move all terms containing a to the left, all other terms to the right. Add '-1abfhor' to each side of the equation. a + -1abfhor = abfhor + -1abfhor + b2fhor Combine like terms: abfhor + -1abfhor = 0 a + -1abfhor = 0 + b2fhor a + -1abfhor = b2fhor Combine like terms: b2fhor + -1b2fhor = 0 a + -1abfhor + -1b2fhor = 0 The solution to this equation could not be determined.

OpenStudy (jordanloveangel):

hop this helped

OpenStudy (anonymous):

yeah im lost but thanks (:

OpenStudy (jordanloveangel):

lol thanx this is the best i can do u still need help and if u do anytime #me kk

OpenStudy (anonymous):

okay (: